OpenStudy (anonymous):

The longer leg of a right triangle is 1 ft longer than the shorter leg. The hypotenuse is 1 ft shorter than twice the shorter leg. Find the length of the shorter leg of the triangle.

OpenStudy (anonymous):

let shorter leg=a longer leg=a+1 hypo=2a-1 a^2+(a+1)^2=(2a-1)^2 2a^2-6a=0 a=3

OpenStudy (anonymous):

shorter leg= x long leg = 1ft+x hyp= 2x-1ft short leg= 3 correct given above